if X/Y=a+2/a-2 Than Proved that  x2-y2/x2+y2=4a/a2+4

Grade:10

1 Answers

Ayush
6 years ago
It can be solved using componendo and dividendo.» x/y = a+2/a-2» x²/y² = a²+4+4a/a²+4-4a» x²-y²/x²+y² = 8a/2a²+4 = 4a/a²+4.Hence proved!😃